OpenStudy (anonymous):

\[(7+\sqrt{2})(8-3\sqrt{2})\]

OpenStudy (anonymous):

okay, so all we have to do is FOIL.

OpenStudy (anonymous):

you should get: \[50-13\sqrt{2}\]

OpenStudy (anonymous):

i did thank you

OpenStudy (anonymous):

Rationalize the denominator of

OpenStudy (anonymous):

\[(7)(8)+(7)(-3\sqrt{2})+(\sqrt{2})(8)+(\sqrt{2})(-3\sqrt{2})\] \[56-21\sqrt{2}+8\sqrt{2}-6\] \[50-13\sqrt{2}\]

OpenStudy (anonymous):

you can click the good answer button to thank people for their answer

OpenStudy (anonymous):

\[\sqrt{50}/\sqrt{5}\]

OpenStudy (anonymous):

how do u rationalize the denominator

OpenStudy (anonymous):

multiply to and bottom of that fraction by \[\sqrt{5}\]

OpenStudy (anonymous):

i am sorry thats wrong

OpenStudy (anonymous):

i got \[\sqrt{10}\]

OpenStudy (anonymous):

yeah, thats right

OpenStudy (anonymous):

i was thinking about something else

OpenStudy (anonymous):

or is it 5\[5\sqrt{5}/2\]

OpenStudy (anonymous):

no, its is \[\sqrt{10}\]

OpenStudy (anonymous):

how about Rationalize the denominator of \[3/\sqrt{2}\]

OpenStudy (anonymous):

When we multiply the top and bottom by \[\sqrt{2}\] we get : \[(3\sqrt{2})/(2)\]